# gen2seg: Generative Models Enable Generalizable Instance Segmentation

### v1

@@ 3.2 Promptable Instance Segmentation
$S_{p}(x,y)=\frac{1}{\|F(x,y)-q_{p}\|_{2}},$ (7)
and normalize it to $[0,1]$, and smooth it with a joint bilateral filter [59, 46] (using $F$ as guidance), thus averaging the similarities close in both pixel location $(x,y)$ and feature value $F(x,y)$. Finally, assuming a set of $k$ point prompts, we merge all corresponding $S_{p}$ by taking the per-pixel maximum across $k$ similarity maps, and threshold the merged similarity map to produce the binary mask.
@@ 4.1 Datasets
Training: Inspired by previous work [28], we train our model exclusively on synthetic data. We combine two datasets: Hypersim [51] and Virtual Kitti 2 [4]. Hypersim provides a rich variety of indoor scenes (i.e. bathrooms, bedrooms, libraries, and kitchens), while Virtual Kitti 2 focuses on outdoor driving scenes, with annotations limited to cars. Both datasets are realistic and do not contain other styles. After removing images with too few masks, our dataset comprises of 86,000 images (66,000 from Hypersim and 20,000 from Virtual Kitti 2), with a sampling strategy that selects Hypersim images 90% of the time. Neither dataset includes annotations of people, animals, and several other categories. Additionally, while the number of images is comparable to existing instance segmentation datasets, the diversity is substantially lower. Our subset of Hypersim contains multiple views sampled from just 457 scenes, while Virtual Kitti 2 contains just 5 videos (each $\sim$15 seconds long), repeated from different viewing angles and weather conditions. A list of labeled object types in Hypersim is available in Appendix D.
Evaluation: Our finetuned models have seen masks of some limited object categories in a single synthetic and realistic style. We aim to evaluate our models’ zero-shot generalization to unseen categories and styles. We select 5 datasets from [29], each of which contains a very different domain from the others: $\text{COCO}_{\text{exc}}$ (COCO [36], but we exclude object types seen in finetuning), DRAM [11] (art), EgoHOS [64] (egocentric), iShape [62] (complex and fine structures), and PIDRay [63] (luggage x-rays). We describe the details and motivation for each dataset in Appendix A.2.
4.2 Models and Baselines
Understanding the conditions under which generalization is possible is central to understanding our method. Thus, we implement our method on several generative models and compare them with some baselines. Unless specified, the models are finetuned using the loss and datasets described above. We finetune Stable Diffusion variants at 480$\times$640 (Hypersim) and 368$\times$1024 (Virtual Kitti 2) and ImageNet-pretrained models at 224$\times$224.
We apply our method to four different settings that include diffusion and MAE models: Stable Diffusion v2 (SD) [52]: Finetuned end-to-end by fixing the timestep to the highest. MAE with Decoder (MAE-B/H) [21]: We finetune an MAE with its decoder end-to-end to showcase that a strong generative prior learned solely from ImageNet-1K images without internet-scale pretraining or text supervision can effectively generalize. We use MAE ViT-B for direct comparison to DINO-B/SimpleClick (see below) and MAE ViT-H to demonstrate scalability. Stable Diffusion v2 - Category Limited (SD-Limited): Finetuned just like SD, but we ignore loss on pixels within the bounding boxes of unknown-category masks to make sure the validation categories are not seen as the unknown category masks at the training time. This demonstrates generalization under a very limited set of categories.
Method $\text{COCO}_{\text{exc}}$-L $\text{COCO}_{\text{exc}}$-M $\text{COCO}_{\text{exc}}$-S DRAM EgoHOS iShape PIDRay
@@ 4.2 Models and Baselines
Table 1: We evaluate zero-shot mIoU at a single prompt point on a wide spread of datasets. We match or recover a high percentage of performance (minimum 70%) on all datasets except $\text{COCO}_{\text{exc}}$-M/S. This suggests that, for larger objects, our models have learned strong object-level representations that transfer across categories and styles. Our baselines, SimpleClick and DINO-B are far below MAE-B, suggesting this generalization is unique to generative models. Additionally, we strongly outperform SAM on the iShape dataset, which evaluates segmentation of detailed and complex structures.
@@
We compare with the following baselines: SimpleClick [39]: A SOTA promptable segmenter using an MAE-B ViTDet [34] backbone. We finetune SimpleClick using its released training code on our synthetic dataset to show that existing architectures cannot generalize well beyond supervised categories. DINO + VAE (DINO-B): To test whether generalizable object groupings are exclusive to generative pretraining, we attach DINO [7] to a frozen VAE (from Stable Diffusion) via a simple up-conv and fine-tune end-to-end. DINO provides the discriminative features, while the VAE can decode to object shapes unseen in finetuning. Segment Anything (SAM) [29]: We use SAM ViT-H off-the-shelf as a high-water mark for generalization that is supervised on the huge SA-1B dataset with 1B annotated masks. Additional details are in Appendix A.1.
4.3 Qualitative Results
We show several qualitative results of our method in Figure 2. These results are selected to highlight potential advantages of generative models over traditional methods. We highlight our model’s ability to represent scenes compositionally in Figure 3. We also showcase masks generated from our features and prompting, compared with our baselines in Figure 4. Additionally, for fair comparison, we randomly sample one page of figures without any cherry picking for each dataset in the appendix (Figures 8, 9, 10, 11, and 12). Finally, we showcase some high resolution results on challenging, in-the-wild images in Figures 13, 14, 15, 16, 17, 18, and 19 (again in the appendix).
Figure 5: We evaluate segmentation quality as the number of prompt points increases. Our SD model marginally exceeds SAM at 1 prompt point and recovers $>$82% of SAM’s performance at 9 prompt points. This is surprising as we do not use a learned mask encoder for multiple prompt points, but simply merge similarity maps computed individually from each point prompt.
Figure 6: We plot a precision-recall of the zero-shot edge detection on BSDS500. Our strongest models outperform SAM’s precision when recall is low, suggesting their segmentations lie on the exact boundaries of the corresponding objects. Our MAE-H model matches SAM’s performance at low recall, despite being trained at only 4% of the resolution ($224\times 224$ vs. $1024\times 1024$).
@@ 4.4 Zero-Shot Promptable Instance Segmentation
We evaluate our model on the task of zero-shot promptable instance segmentation. Following [29], we first evaluate our models’ ability to produce reasonable masks using only a single prompt point at the ground truth object center and comparing the predicted mask’s IoU with the respective ground truth. Then, following the so-called "golden" standard of prompting [49, 29, 39, 55, 38, 54], we iteratively find the largest contiguous area of the ground truth with no mask predicted yet, and select the next prompt point in that area closest to the area’s center. As promptable segmentation is a highly ambiguous task, we do not necessarily expect to get high IoU, but we do hope for it to approach our high-water mark, SAM.
@@
Results. First, we examine model performance using a single prompt at the object center (Table 1). Except for $\text{COCO}_{\text{exc}}$-M/S (Medium and Small), our model approaches or marginally exceeds SAM’s performance, despite never seeing labeled masks for these categories. This indicates that generative models learn transferable object-level features, particularly for larger objects and intricate details (evidenced by superior results on the iShape dataset). Interestingly, SAM struggles in some cases because it seems to group by texture for out-of-distribution cases, such as art. Our SD-Limited model is able to generalize although unknown masks (over 30% of the dataset) are removed, but sometimes struggles to discriminate visually similar adjacent objects (as seen in Figure 13).
Our method also surpasses both baselines, SimpleClick and DINO-B. SimpleClick, as expected, struggles with zero-shot mask generation since its feature pyramid, trained from scratch on our synthetic datasets with limited object types, lacks relevant priors.
@@
Our finetuned DINO-B model successfully activates on objects, but struggles to separate their instances. We hypothesize this is because self-distillation (and discriminative pretraining in general) over-emphasize semantics via invariant representations, meaning they enforce that the output representation does not change across augmentations. For example, discriminative pretraining forces the model to learn the same representation for both an image with two cats and its crop with only a single cat. In contrast, to succeed at instance segmentation, one must learn equivariant representations, meaning they account for changes in the scale, shape, or structure of the image (and the objects within) [18]. We hypothesize that generative models are well posed to learn equivariant representations because they must learn to synthesize a plausible image for every corrupted input they receive.
However, our models have limitations segmenting small objects, likely due to biases from pretraining: Stable Diffusion’s text conditioning emphasizes large, prominent objects, while MAE’s ImageNet-1K pretraining prefers central "main" objects. Additionally, models like SAM and other state-of-the-art segmenters [19, 10, 5] fine-tune at 1024$\times$1024 resolution, whereas we fine-tune at lower resolutions: 480$\times$640 (Hypersim) and 368$\times$1024 (Virtual Kitti 2) for Stable Diffusion variants, and 224$\times$224 for ImageNet-based models. We expect that stronger generative models, such as FLUX.1, would enhance small-object segmentation. We leave this to future work due to their large parameter counts.
Additionally, we explore if our masks improve with additional prompt points (Figure 6). We average IoU across datasets, excluding $\text{COCO}_{\text{exc}}$-M/S (since few small objects are detected in the first place). Remarkably, our SD model achieves over 82% IoU relative to SAM at 9 prompt points, despite lacking a learned prompt encoder for multi-point prompts or masks from any of the evaluated categories.

### v2

@@ 3.2 Point-Promptable Instance Segmentation
$S_{p}(x,y)=\min(1,\frac{1}{\|F(x,y)-q_{p}\|_{2}}),$ (7)
normalize it between [0, 1], and smooth it with a joint bilateral filter (Petschnigg et al., 2004) (using $F$ as guidance), thus averaging the similarities close in both pixel location $(x,y)$ and feature value $F(x,y)$. Finally, assuming a set of $k$ point prompts, we take the per‐pixel maximum across $k$ similarity maps, and threshold the merged similarity map to produce the binary mask.
@@ 4.1 Datasets
Method $\text{COCO}_{\text{exc}}^{L}$ $\text{COCO}_{\text{exc}}^{M}$ $\text{COCO}_{\text{exc}}^{S}$ DRAM EgoHOS iShape PIDRay
@@
Table 1: We evaluate zero-shot mIoU at a single prompt point on a wide spread of datasets. We match or recover a high percentage of performance (minimum 70%) on all datasets except $\text{COCO}_{\text{exc}}^{M}$/S. This suggests that, for larger objects, our models have learned strong object-level representations that transfer across categories and styles. Our baselines, SimpleClick and DINO-B are far below MAE-B, suggesting this generalization is unique to generative models. Additionally, we strongly outperform SAM on the iShape dataset, which evaluates segmentation of detailed and complex structures. SAM is trained on SA-1B. All other models are trained on our limited Hypersim+VK2.
Training: Inspired by previous work (Ke et al., 2024), we train our model exclusively on synthetic data. We combine two datasets: Hypersim (Roberts et al., 2021) and Virtual Kitti 2 (Cabon et al., 2020). Hypersim provides a rich variety of indoor scenes (i.e. bathrooms, bedrooms, libraries, and kitchens), while Virtual Kitti 2 focuses on outdoor driving scenes, with annotations limited to cars. Both datasets are realistic and do not contain other styles. After removing images with too few masks, our dataset comprises of 86,000 images (66,000 from Hypersim and 20,000 from Virtual Kitti 2), with a sampling strategy that selects Hypersim images 90% of the time. Neither dataset includes annotations of people, animals, and several other categories. Additionally, while the number of images is comparable to existing instance segmentation datasets, the diversity is substantially lower. Our subset of Hypersim contains multiple views sampled from just 457 scenes, while Virtual Kitti 2 contains just 5 videos (each $\sim$15 seconds long), repeated from different viewing angles and weather conditions. A list of labeled object types in Hypersim is available in Appendix D.
Evaluation: Our finetuned models have seen masks of some limited object categories in a single synthetic and realistic style. We aim to evaluate our models’ zero-shot generalization to unseen categories and styles. We select 5 datasets from (Kirillov et al., 2023), each of which contains a very different domain from the others: $\text{COCO}_{\text{exc}}$ (The COCO 2017 validation set (Lin et al., 2014), but we exclude object types seen in finetuning, The list of categories are presented in Appendix D), DRAM (Cohen et al., 2022) (art), EgoHOS (Zhang et al., 2022b) (egocentric), iShape (Yang et al., 2021) (complex and fine structures), and PIDRay (Zhang et al., 2022a) (luggage x-rays). We describe the details and motivation for each dataset in Appendix A.2.
4.2 Models and Baselines
[…]
We apply our method to four different settings that include diffusion and MAE models: Stable Diffusion v2 (SD) (Rombach et al., 2021): Following (Garcia et al., 2024), we set t=N1t=N-1 in an $N$-step latent diffusion model and finetune end-to-end without noise. Using the frozen VAE to decode the U-Net output, we optimize pixel-space loss, yielding a one-step, deterministic image segmenter. MAE with Decoder (MAE-B/H) (He et al., 2022): We finetune an MAE with its decoder end-to-end to showcase that a strong generative prior learned solely from ImageNet-1K images without internet-scale pretraining or text supervision can effectively generalize. We use MAE ViT-B for direct comparison to DINO-B/SimpleClick (see below) and MAE ViT-H to demonstrate scalability.
@@ 4.2 Models and Baselines
We compare with the following baselines: SimpleClick (Liu et al., 2023a): A SOTA promptable segmenter using an MAE-B ViTDet (Li et al., 2022) backbone. We finetune SimpleClick using its released training code on our synthetic dataset to show that existing architectures cannot generalize well beyond supervised categories. DINO + VAE (DINO-B) (Caron et al., 2021): To test whether generalizable object groupings are exclusive to generative pretraining, we attach DINO to a frozen VAE decoder (from Stable Diffusion) via a simple up-conv and fine-tune end-to-end. DINO provides the discriminative features, while the VAE can decode to object shapes unseen in finetuning. Segment Anything (SAM) (Kirillov et al., 2023): We use SAM ViT-H off-the-shelf as a high-water mark for generalization that is supervised on the huge SA-1B dataset with 1B annotated masks.
Data $\text{COCO}_{\text{exc}}^{L}$ $\text{COCO}_{\text{exc}}^{M}$ $\text{COCO}_{\text{exc}}^{S}$ DRAM EgoHOS iShape PIDRay
Original 50.0/57.6 23.2/38.8 3.5/8.5 40.3/48.2 31.9/40.0 34.9/51.4 24.1/30.9
COCO 53.6/64.0 18.8/44.0 2.9/9.8 48.1/51.2 26.8/35.2 33.4/41.2 25.7/31.9
ClevrTex 40.0/47.1 19.4/21.6 1.9/7.5 23.5/28.0 21.4/21.9 27.6/32.1 22.2/23.7
10 classes 54.8/56.1 21.7/35.2 2.7/5.0 40.1/45.1 29.4/38.5 33.0/53.6 17.6/22.8
5 classes 42.1/47.6 16.2/29.7 1.4/6.4 34.2/38.2 23.7/34.4 28.5/48.5 15.2/19.4
Table 2: We explore how the diversity of the finetuning domain impacts generalization. Interestingly, our model still performs well with real-world data (COCO) or synthetic shape-segmentation datasets (ClevrTex). Furthermore, performance with just 10 classes from Hypersim results in nearly identical performance to the full dataset (33+ classes), suggesting generalization emerges without diverse finetuning data. However, some degree of diversity is still needed for optimal performance, as seen by the performance drops on ClevrTex or with only 5 classes. We report results with both the MAE-H and SD backbones as MAE-H/SD.
We examine the role of training data by varying both the data domain and number of object categories. We train on (i) the full MS-COCO train split or (ii) ClevrTex (Karazija et al., 2021), a synthetic, shape-centric dataset. COCO has real-world visual diversity but includes noisier labels while ClevrTex contains pixel-perfect annotations but a narrow set of simple object types. As our Hypersim and Virtual Kitti 2 mix contains both complex shapes and high quality labels, comparing these regimes helps disentangle the impact of object complexity from annotation quality. To study category diversity, we also experiment with restricting Hypersim to only 10 and 5 classes of labeled masks in finetuning. To ensure the model does not see the masks of unknown categories, we disable the loss for pixels within the bounding box of all unknown objects. Additional details regarding models and datasets appear in Appendix A.1.
@@ 4.3 Zero-Shot Promptable Instance Segmentation
We evaluate our model on the task of zero-shot point-promptable instance segmentation. Following Kirillov et al. (2023), we first evaluate our models’ ability to produce reasonable masks using only a single prompt point at the ground truth object center and comparing the predicted mask’s IoU with the respective ground truth. Then, following the so-called “golden” standard of prompting (Ravi et al., 2024; Kirillov et al., 2023; Liu et al., 2023a; Sofiiuk et al., 2022; Lin et al., 2022; Sofiiuk et al., 2020), we iteratively find the largest contiguous area of the ground truth with no mask predicted yet, and select the next prompt point in that area closest to the area’s center. As promptable segmentation is a highly ambiguous task, we do not necessarily expect to get high IoU, but we do hope for it to approach our high-water mark, SAM.
@@
Results. First, we examine model performance using a single prompt at the object center (Table 1). On all evaluation datasets except $\text{COCO}_{\text{exc}}^{M/S}$ (Medium and Small), our model approaches or marginally exceeds SAM, despite never seeing labeled masks for these categories. This indicates that generative models learn transferable object-level features, particularly for larger objects and intricate details (evidenced by superior results on the iShape dataset). Interestingly, SAM struggles in some cases because it seems to group by texture for out-of-distribution cases, such as art.
Our method also surpasses both baselines, SimpleClick and DINO-B. SimpleClick, as expected, struggles with zero-shot mask generation. SimpleClick’s failure to generalize represents a weakness in existing segmentation architectures: nearly all models use mask predictors finetuned from scratch (Kirillov et al., 2023; Wang et al., 2024). When faced with object types unlike anything seen in finetuning, these models will fail to generalize as the mask predictor has only seen the object types in finetuning and lacks any other visual priors. Our method deliberately uses only generatively pretrained parameters, so that the entire model retains broad visual priors for test-time generalization.
@@
Our finetuned DINO-B model successfully activates on objects, but struggles to separate their instances. We hypothesize this is because self-distillation (and discriminative pretraining in general) over-emphasize semantics via invariant representations, meaning they enforce that the output representation does not change across augmentations. In contrast, to succeed at instance segmentation, one must learn equivariant representations, meaning they account for changes in the scale, shape, or structure of the image (and the objects within) (He, 2017). We hypothesize that generative models are well posed to learn equivariant representations because they must learn to synthesize a plausible image for every corrupted input they receive.
[…]
We also find that our models generalize well, even without synthetic or complex datasets, as shown in Table 2. Surprisingly, our models learn to generalize even with only 5 object types (books, chairs, lamps, tables, and pillows) seen in finetuning, or when only finetuned on simple shapes such as cubes or spheres (ClevrTex). Furthermore, our models still excel at segmenting fine structures (as shown on iShape) when finetuned on COCO, which mostly contains polygonal mask annotations. Additionally, we see only minor improvement when generative models are finetuned with COCO, suggesting that our models’ zero-shot generalization is very close to the “upper bound” of seeing the objects in finetuning. This suggests generative pretraining is a powerful paradigm to learn generalizable instance grouping priors.
SAM DINO-B MAE-B MAE-H SD MAE-H (COCO) SD (COCO) MAE-H (ClevrTex)
79.0 33.2 53.7 80.5 93.4 75.9 89.7 86.4
SD (ClevrTex) MAE-H (10 classes) SD (10 classes) MAE-H(5 Classes) SD (5 Classes) Sobel
88.5 86.9 92.5 83.1 91.7 65.0
Table 3: We evaluate the zero-shot edge AP for recall less than 20% of the edges synthesized from our models and baseline segmentations on BSDS500. Nearly all generative models outperform SAM, even when trained on polygonal mask edges from COCO. This suggests that our models’ ability to produce detailed groupings is due to the generative prior, and not any specific dataset.
4.4 Zero-Shot Edge Detection
We evaluate our segmentation features on the task of edge detection with the BSDS500 dataset (Martin et al., 2001). It is important to note that we are not simply trying to find all edges in the image, but only object boundaries. Following Kirillov et al. (2023), we simply apply a Sobel filter on the predicted features followed by nonmax suppression to thin the edges. We use the edge detection described above on the original image as a “weak” baseline and on the output of SAM’s AutoMaskGenerator as a “strong” baseline. We report AP for recall less than 20%. We explain this choice (and display the full precision-recall curves) in Appendix B.
Results. Our model accurately delineates the edges of primary objects in each image despite never seeing a mask of the object type. As shown in Table 3, our SD model produces much finer edges compared to SAM. Additionally, our MAE-H model marginally outperforms SAM despite being supervised at less than 5% of the resolution (224 $\times$ 224 vs 1024 $\times$ 1024). We also observe that when we change the training data from synthetic Hypersim+VK2 to human-labeled COCO, we observe a decrease of less than 5 points in edge AP on both SD and MAE-H models. Because COCO’s edges are coarse and polygonal, the relatively small performance drop supports our hypothesis that our model’s accurate and clean edges stem from generative pre-training and are not due to the bias from the synthetic data. More interestingly, our SD (COCO) model is still better than SAM by more than 10 points for recall less than 20%. Both COCO and SA-1B have similar polygonal edges, which suggests by learning to synthesize scenes, generative models inherently learn a detailed representation of object boundaries. Our precise edges also persist (and are sometimes even stronger than the base model) when dataset complexity reduces, suggesting the generative model “defaults” to predicting clean edges when segmenting an object.
5 Conclusion
Our findings suggest that learning to generate visual reality inherently teaches a detailed understanding of its constituent parts. Our models are able to segment objects and styles nothing like anything any labels seen in finetuning, yet still achieve competitive performance with heavily supervised models. As we continue to scale generative models and diversify their pretraining data, their ability to perceive the visual world will only grow. Learning to leverage these powerful representations for a wide variety of visual tasks has the potential to enable a new frontier in generalizable perception.
Appendix
Appendix A Experimental Models and Datasets
A.1 Additional Details for Models and Baselines
Stable Diffusion 2 (SD): We finetune Stable Diffusion 2 U-Net end-to-end using our loss. We follow the method in (Garcia et al., 2024) which enables deterministic one-step prediction. This has been shown to outperform multi-step stochastic inference with standard diffusion training for other perceptual tasks. To train it in pixel space, we fix the timestep to the highest (999). We replace the input with our image’s VAE latent without adding any noise. We set the CLIP embedding to the null condition. We freeze the VAE and finetune the U-net end to end with our instance coloring loss as described above. We show in Appendix B that the choice of timestep has negligible effect on the results as the timestep cross-attention within the U-net is finetuned too.
DINO + VAE (DINO-B): We investigate whether a strong discriminative model can suffice for segmentation features without necessarily learning to synthesize images. However, we must pair it with a model that knows how to upsample images of all types from a low-dimensional space so that it will be able to generate images of objects unseen in training. To investigate this, we join ImageNet-1K pretrained DINO ViT-B with the Stable Diffusion VAE decoder, and connect them with a single up-conv layer. Additionally, the latent features of images compressed with the Stable Diffusion VAE often look like the image itself (Kouzelis et al., 2025). Thus, to succeed, all DINO needs to do is generate these object groupings at a lower resolution.
Segment Anything (SAM): SAM is a large-scale promptable instance segmentation model supervised on over a billion masks from a very large distribution of data. We use SAM as a benchmark to evaluate the extent to which our model generalizes across a wide variety of images. While in some cases we outperform SAM, we do not mean for this to claim that our models are inherently superior to it. We use the publicly available ViT-H checkpoint and inference library.
Dataset Baselines: We finetune the models (MAE-H and SD) with the same loss and implementation details. For 5 and 10 category models, we computed no loss for any pixels inside the bounding boxes of the objects that are outside the “allowed” categories. We also pruned any images with more than 35% of image area that is “not allowed” or any images with no allowed masks as we found it important to have a smaller amount of high quality data, rather than a larger set of low quality data. Because of limited time, we reported results for SD (10 classes) and SD (5 classes) after only 15k iterations of finetuning, rather than the full 30k. We will update our rebuttal draft with results from the full finetuning schedule.
To match their pretraining resolutions, SimpleClick, DINO, and MAE are finetuned at $\times224$ resolution, rather than the standard ones described in implementation details.
A.2 Evaluation Datasets
PIDRay: Humans effortlessly apply their understanding of object shapes even in scenarios never encountered in nature. For instance, TSA agents can quickly identify dangerous items in X-ray images of luggage. To assess our models’ generalization to this challenging scenario, we employ the PIDRay dataset, which features labeled examples of hazardous objects in baggage. This task is particularly demanding because many dangerous items are small and deliberately concealed within other objects. We evaluate on the “easy” subset of the test set.
Appendix B Additional Results and Experimental Details
Figure 5: We evaluate segmentation quality as the number of prompt points increases. Our SD model marginally exceeds SAM at 1 prompt point and recovers >>82% of SAM’s performance at 9 prompt points. This is surprising as we do not use a learned mask encoder for multiple prompt points, but simply merge similarity maps computed individually from each point prompt.
Figure 6: We plot the full precision-recall of the zero-shot edge detection on BSDS500. Our strongest models outperform SAM’s precision when recall is low, suggesting their segmentations lie on the exact boundaries of the corresponding objects. However, as shown in Figure 7, our models sometimes do not predict objects for certain regions of the image, leading to lower precision at higher recall values.